OpenStudy (anonymous):

Suppose 'f' is a function with domain [-13,17] and range (2,7]. Find the domain and range of both f(5x+2) and 5 f(x) +2

OpenStudy (anonymous):

@genius12

OpenStudy (zzr0ck3r):

-13<=5x+2<=17 -15<=5x<=15 -3<=x<=3 so domain is [-3,3] for range multiply every number in the current range by 5 then add 2 (12,37]

OpenStudy (anonymous):

range is incorrect

OpenStudy (zzr0ck3r):

do you know the answer? I don't see how that is wrong we have range given by f(x) and we need range given by 5f(x) + 2

OpenStudy (anonymous):

i believe your correct, however what is the domain for 5 f(x) +2

OpenStudy (zzr0ck3r):

the same as f(x)

OpenStudy (anonymous):

cheers